### v2

# Shot noise in Aharonov-Bohm interferometers: Comparison of helical and conventional setups

Abstract
We study tunneling transport through quantum Aharonov-Bohm (AB) interferometers and demonstrate that interference effects strongly modify shot noise of the current. We discuss in detail two simplest setups: conventional single-channel spinless interferometer and interferometer formed by helical edge states of two-dimensional topological insulator. We demonstrate that both in the conventional and the helical case the interference dramatically changes the Fano factor and its magnetic field dependence. For weak tunneling coupling, the Fano factor of both setups exhibits a periodic series of sharp AB peaks depending on the magnetic flux piercing the system. Our key finding is that the Fano factor in the helical interferometer provides information about the presence of backscattering defects violating topological protection. In particular, the amplitude of AB peaks in the helical setup is proportional to the strength of the defect in contrast to conventional setup, where peaks have finite amplitude even in the ballistic case.
I Introduction
Quantum interferometry is a rapidly growing area of electronics with a huge potential for practical use, in particular, in the booming fields of quantum computing and quantum cryptography. One of the beautiful manifestation of quantum interference is the Aharonov-Bohm effect Aharonov and Bohm (1959, 1963)—the sensitivity of the phase of electron wave function to magnetic flux. Quantum interferometers based on this effect are used for high-precision measurements of magnetic fields, for example, by using SQUID circuits.
The simplest setup to study interference and particularly the Aharonov-Bohm effect is a ring-shaped interferometer with two single-channel arms placed in the perpendicular magnetic field. Due to the interference of electron trajectories winding around such a setup, the observables, like transmission coefficient and the noise intensity, show Aharonov-Bohm oscillations with a certain periods $\Delta\phi$ as a function of dimensionless magnetic flux $\phi=\Phi/\Phi_{0}$ threading the device, where $\Phi_{0}=hc/e$ is the flux quantum and $\Phi=BS,$ where $B$ is homogeneous magnetic field and $S$ is the area of the region encompassed by one-dimensional channel. The key statement of our paper is that noise in such a simple system is still not well understood and depends on specific realizations of the single-channel arms.
Here, we will investigate and compare in details two types of setups: (i) a spinless single-channel interferometer based on conventional materials, i.e. a single-channel spinless quantum wire, folded in a ring-like shape and having two tunnel or metallic contacts attached (see Fig. 1a); (ii) an interferometer based on helical edge states. Below we refer to these systems as a conventional interferometer (CI) and a helical interferometer (HI), respectively. The latter interferometer can be implemented in two-dimensional topological insulators  Bernevig and Hughes (2013), which are insulators in the bulk, but have helical one-dimensional (1D) edge states conducting the current without dissipation Bernevig and Hughes (2013); Hasan and Kane (2010); Qi and Zhang (2011). These edge states are formed by electrons with opposite spins traveling in opposite directions, so that backscattering by non-magnetic impurities is prohibited. If one attaches two contacts to the edge and apply the gate voltage to shift the Fermi level into the band gap, then the conductance of such a device is fully determined by the edge states (see Fig. 1b). Since the sample boundary can be bypassed in two directions, possibly with several windings, such a system represents an interferometer that encodes information about properties of the helical edge states (HES).
Both setups, CI and HI, have two electron waves propagating clockwise and counterclockwise in 1D channel. However, the interference phenomena in these systems manifest themselves in a rather different way.
A spinless CI based on conventional materials (Fig. 1a) shows interference effects even in the simplest ballistic case, when both arms of interferometer do not contain any impurities. By contrast, the waves, propagating in opposite directions in ballistic HI, have also opposite spin projections and cannot interfere provided that contacts to the HES are non-magnetic (note that CI differs from Rashba-interferometers, where ballistic interference is possible due to the Aharonov-Casher effect Aronov and Lyanda-Geller (1993); Koga et al. (2004); Nitta and Bergsten (2007); Shmakov et al. (2012), see also Ref. Manchon et al. (2015) for review and references therein). Interference effects in HI appear in presence of at least one backscattering center (see Fig. 1b). It means that the study of interference effects can help to solve the well-known problem of the lack of topological protection in real samples (see recent experiment Olshanetsky et al. (2023) and discussion of the experimental situation therein). Specifically, observation of interference phenomena in the HI can give information about mechanisms of possible violation of the topological protection.
Figure 1: Aharonov-Bohm interferometers based on conventional spinless single-channel quantum wire (a) and on helical edge channels of 2D topological insulator (b). The leads (shown by gray color) are modeled by conventional single-channel [spinless (a) and spinful (b)] wires. The black dot in panel (b) represents a backscattering impurity. Gray regions represent contacts described by $3\times 3$ and $4\times 4$ $S$-matrix for CI and HI, respectively [see Eqs. (2), (4) and Figs. 2, 4]. Homogeneous magnetic field, $B,$ is perpendicular to the picture plane. Corresponding magnetic flux, $\Phi=BS,$ is proportional to the area, $S,$ of the region encompassed by one-dimensional channels
For both setups we will investigate the manifestations of quantum interference in the shot noise of the current. Such noise is a consequence of the discreteness of the electron charge. Importantly, its measurement provides information not available from conductance measurements: namely, the charge and statistics of current carriers, and the internal energy scales of the system de Jong and Beenakker (1997); Blanter and Büttiker (2000). Our primary goal is to calculate the Fano factor, $\mathcal{F}$, which is the ratio of the shot noise and the so-called Schottky noise value. This ratio is often used to characterize different transport regimes. For example, the values $\mathcal{F}=1/3$~ and $\mathcal{F}=1/2$~ for a diffusive conductor and tunneling transmission through a quantum level, respectively, are universal and independent of the system details de Jong and Beenakker (1997); Blanter and Büttiker (2000). As we demonstrate below, the Fano factor also is very sensitive to the interference effects which manifest itself in different way in CI and HI.
It is worth noting that the shot noise in HES has already been discussed for the case of infinite edge Lezmy et al. (2012); Del Maestro et al. (2013); Tikhonov et al. (2015); Mani and Benjamin (2017); Väyrynen and Glazman (2017); Piatrusha et al. (2018); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022); Mishra and Benjamin (2023), i.e. without taking into account the interference effects. In particular, a great attention was paid to infinite HES with a dynamic magnetic impurity Väyrynen and Glazman (2017); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022)), whose magnetic moment changes direction after each scattering event. Discussing HI in this paper, we focus on the opposite case of static backscattering defect, which is not necessarily magnetic (in particular, backscattering mechanisms involving charged puddles in the bulk of topological insulator are actively discussed Väyrynen et al. (2013)). We will describe such defect by $S$ matrix of general type. The experimental measurement of the Fano factor of HI at zero magnetic field for the edge states of 2D TI gives the value $0.1<\mathcal{F}<0.3$  Tikhonov et al. (2015); Piatrusha et al. (2018). The upper value, 0.3, is close to the value $1/3$ for diffusive conductor. A similar result was obtained in the model with a large number of “islands”, allowing spin relaxation and tunnel-connected to the HES Aseev and Nagaev (2016). The effect of islands of different types on HES is currently being actively debated (see also recent works Olshanetsky et al. (2023); Krainov et al. (2025) and references therein).
Since the conductance and shot noise of interferometers of various types were also studied in a great number of works, we will start with a short description of the state of the art. Let us first recall the key results known for the conductance of the Aharonov-Bohm interferometer. At weak tunnel coupling and at low temperatures, the conductance $G(\phi)$ exhibits sharp resonances at certain values of magnetic flux, arising when quantum levels of the system are crossed by the Fermi level Büttiker et al. (1984, 1985).
As was understood later Jagla and Balseiro (1993), the case of relatively high temperatures, $T$, is much less trivial. One might suggest that interference is suppressed when $T$ becomes larger than the level spacing, $\Delta.$ However, as was first predicted in Ref. Jagla and Balseiro (1993) for CI, the conductance shows interesting interference-induced behavior in this case. Specifically, as it was shown theoretically, for both CI Jagla and Balseiro (1993); Dmitriev et al. (2010); Shmakov et al. (2012, 2013); Dmitriev et al. (2015, 2017) and HI Niyazov et al. (2018, 2020, 2021a, 2021b), the interference effects lead to resonance behavior of the conductance $G$ even in the case
$T\gg\Delta=2\pi v_{\rm F}/L,$ (1)
where $L=L_{1}+L_{2},$ with $L_{1,2}$ the lengths of the interferometer arms, and $v_{\rm F}$ is the Fermi velocity. However, instead of resonances, $G(\phi)$ exhibits narrow antiresonances under the condition (1). Physically, anti-resonances arise due to trajectories that interfere destructively at any energy $\varepsilon$ and, accordingly, are insensitive to energy averaging (see discussions in Dmitriev et al. (2010) and Niyazov et al. (2018) regarding CI and HI, respectively).
The conductance in CI in the temperature range (1) has been studied in detail, in particular, considering the effects of the electron-electron interaction Dmitriev et al. (2010), disorder Shmakov et al. (2013), and spin-orbit interaction Shmakov et al. (2012). Recently, some of these results were generalized to HI systems: to single ring-shaped HI Niyazov et al. (2018, 2020, 2021a, 2021b) with a static magnetic defect and, partially, to arrays of coupled helical rings Niyazov et al. (2023). (See also discussion for other types of interferometers Feldman et al. (2007); Wang and Feldman (2010); Yang (2015).)
Surprisingly, non-trivial interference-induced characteristics of noise emerging under condition (1) were not discussed previously even for the simplest possible model of a single-channel CI, in view of the huge number of publications devoted to CI. To the best of our knowledge, the influence of the interference effects on the noise has previously been discussed only in the low-temperature regime, $T\ll\Delta$, see review in Ref. Kobayashi and Hashisaka (2021) for CI and Refs.  Edge et al. (2013); Dolcini (2015) for HI.
On the other hand there is a clear need of the corresponding theory both for CI and HI, which would allow direct comparison with experiment. Indeed, nanoscale rings made from conventional materials with a few or single conducting channels have been already fabricated Shea et al. (2000); Piazza et al. (2000); Fuhrer et al. (2001); Keyser et al. (2003); Zou et al. (2007). Also, the experiment aimed to study HI was recently reported Munyan et al. (2023), although it focused on the low-temperature case. Note that in the latter experiment a quantum point contacts to the HES have been realized, thus allowing experimental verification of interference effects in conductance and noise of tunneling HI, especially having in mind that the condition (1) does not require very strict restrictions on temperature. Indeed, for typical $v_{\rm F}$ of order of $10^{7}$~cm/s and typical system sizes (>1>1 micron), the value of $\Delta$ does not exceed several degrees Kelvin. This means that the interference effects can be studied at relatively high temperatures, which are relevant for various applications. A comment about magnetic-field-induced phenomena is in order to do. Actually, the magnetic field causes two effects: (i) the AB phase appears and (ii) the topological protection is destroyed. Simple estimates show that the flux quantum, $\Phi\sim\Phi_{0}=hc/e,$ is achieved for sample with micron size in fields $B\sim 3$ Oe, well below the expected magnitude of the fields destroying the edge states Du et al. (2015); Zhang et al. (2014); Hu et al. (2016).
Recently we discussed shot noise in HI Niyazov et al. (2024). We considered HI with identical contacts and calculated the noise power in the regime (1) paying attention to the interference effects. We found the Fano factor strongly depending on the tunneling coupling and strength of backscattering defect.
Here, we further develop the results obtained in Ref. Niyazov et al. (2024). We present a detailed study of the sharp Aharonov-Bohm peaks in the Fano factor of the noise. We also demonstrate that the simultaneous measurement of the conductance and the Fano factor of HI allows one to directly determine the strength of the backscattering defect even in the absence of detailed information about scattering amplitudes characterizing the contacts. Having also in mind the recent experiment Munyan et al. (2023), we generalize here the calculation of the Fano factor of HI for experimentally relevant case of non-equivalent contacts.
In Ref. Niyazov et al. (2024), we also briefly discussed the case of CI, announcing without derivation the formula for the Fano factor of the CI with symmetric arms and identical contacts. In this paper we provide detailed calculations, and generalize our analysis for the experimentally relevant case of asymmetric arms and non-identical contacts Shea et al. (2000); Piazza et al. (2000); Fuhrer et al. (2001); Keyser et al. (2003); Zou et al. (2007). We study in detail the dependence of the noise of a CI on the magnetic field. It turns out that the interference effects are much more pronounced for CI and, consequently, the influence of the magnetic field on the Fano factor is much stronger as compared to HI. The reason is that the interference comes into play for CI already in the ballistic case, when impurities in both arms of CI are absent, while the Aharonov-Bohm peaks in HI are proportional to the backscattering probability.
We will discuss ballistic CI without impurities and defects and also HI with a single static defect (which can be magnetic or non-magnetic) placed into one of the arm of the interferometer. The goal of the work is to find the dependence of Fano factor, $\mathcal{F}$, on tunneling coupling, on the magnetic flux, and, in the case of HI, on the strength of backscattering defect. We will find general analytical expressions for $\mathcal{F}^{\rm CI}$, $\mathcal{F}^{\rm HI}$ and analyze simple limiting cases assuming that inequality (1) is satisfied.
II Model
We calculate the current shot noise in the Aharonov-Bohm (AB) interferometers of two types, CI and HI, depicted in Fig. 1a and Fig. 1b, respectively. They are based on spinless single-channel quantum wire and helical edge states. We model leads by conventional single-channel wires (spinless or spinful for the conventional and helical case, respectively). Although this model of the leads is very simplified, it is commonly used in quantum interferometry, starting from the work Büttiker et al. (1984), since, as we demonstrate below, it allows describing the transition from the metallic leads to the tunnel ones. In particular, this model qualitatively describes quantum point contacts to the interferometer.
II.1 Conventional AB interferometer
The unitarity of scattering matrix of the contact results in backscattering on the contacts in case of CI Dmitriev et al. (2010). Here, we use this scattering matrix with real-valued elements:
$\hat{S}^{\mathrm{CI}}=$ $\begin{pmatrix}t_{\mathrm{r}}&t_{\mathrm{out}}&t_{\mathrm{out}}\\ t_{\mathrm{in}}&t_{\mathrm{b}}&t\\ t_{\mathrm{in}}&t&t_{\mathrm{b}}\\ \end{pmatrix}\,$, (2)
$t=$ $\frac{1}{1+\gamma}\,,\quad t_{\mathrm{b}}=-\frac{\gamma}{1+\gamma}\,$,
$t_{\mathrm{in}}=$ $t_{\mathrm{out}}=\frac{\sqrt{2\gamma}}{1+\gamma}\,,\quad t_{\mathrm{r}}=-\frac{1-\gamma}{1+\gamma}\,.$
More general case of $\hat{S}^{\mathrm{CI}}$ is discussed in Refs. Aristov and Wölfle (2011) and Aristov et al. (2010). The scattering amplitudes entering Eq. (2) obey $2t_{\rm in}^{2}+t_{\rm r}^{2}=t^{2}+t_{\rm b}^{2}+t_{\rm out}^{2}=1.$ The meaning of these amplitudes is illustrated in Fig. 2: $t_{\rm in}$ is the amplitude to enter the interferometer into one of the counter-propagating channels, $t_{\rm out}$ is the amplitude to exit the interferometer, $t$ and $t_{\rm b}$ are respectively the amplitudes of forward and backward scattering on the contact inside the interferometer, and $t_{\rm r}$ is the amplitude of backward scattering for electrons coming from the lead. All scattering amplitudes are parameterized by a single parameter $\gamma\in(0,\infty)$.
Figure 2: Scattering amplitudes entering scattering matrix Eq. (2) for CI. The lead is modeled by spinless 1D wire. For tunneling contact, $|t_{\rm r}|\approx 1,$ while the case of metallic contact corresponds to $t_{\rm r}=0.$ Contact region with area $S_{0}$ is shown by gray color. We assume that $BS_{0}\ll\Phi_{0},$ so that the magnetic field does not affect S-matrix.
The meaning of the parameter $\gamma$ is further illustrated in Fig.  3, see also discussion in Refs. Dmitriev et al. (2010); Aristov et al. (2010, 2017). The case $\gamma\ll 1$ corresponds to tunneling contact, where almost closed interferometer is weakly coupled to the leads, with $\gamma$ being a tunneling transparency (see Fig. 3a). Interestingly, the case $\gamma\to\infty$ also corresponds to weak tunneling coupling. However, this case is strongly different from the case $\gamma\to 0,$ as illustrated in Fig. 3b. Metallic contact is modeled by $\gamma\approx 1$ (see Fig. 3c). The scattering matrices describing these three cases are given by (see Eq. (2)):
$\hat{S}^{\mathrm{CI}}(\gamma=0)=$ $\begin{pmatrix}-1&0&0\\ 0&0&1\\ 0&1&0\\ \end{pmatrix}\,,\,\hat{S}^{\mathrm{CI}}(\gamma=\infty)=$ $\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\\ \end{pmatrix}\,$, (3)
$\hat{S}^{\mathrm{CI}}(\gamma=1)=$ $\begin{pmatrix}0&\sqrt{2}&\sqrt{2}\\ \sqrt{2}&-1&1\\ \sqrt{2}&1&-1\\ \end{pmatrix}/2\,.$
It is worth noting that $t_{b}\neq 0$ for any non-zero $\gamma.$ Below we show that two configurations, Fig. 3a and Fig. 3b, may correspond to the same conductance but to different noise.
Figure 3: Schematic illustration of different types of leads (contact area is shown by gray color): panels (a) and (b) shows two different types of tunneling coupling, corresponding to $\gamma=0$ and $\gamma=\infty,$ respectively; panel (c) illustrates “metallic” contact, $\gamma\approx 1.$ Arrows in panels (a) and (b) show processes with non-zero amplitudes (compare with Fig. 2). Panel (c) corresponds to a “metallic contact”, so that all processes in Fig. 2 are allowed. Scattering matrices, describing cases (a), (b), and (c) are given by Eq. (3).
II.2 Helical AB interferometer
Next we discuss model of HI. We consider 2D TI, assuming that the Fermi level is located in the bulk gap, so that the transport between two contacts connected to the system is completely determined by the HES.
Figure 4: The scattering amplitudes entering scattering matrix of contacts, Eq. (4), which are modeled as a single-channel spinful wire. It is assumed that there is no spin flip at the contact and the two spin polarizations at the contact (shown by red and blue color) are completely separated. The $t\to 1$ case corresponds to a tunnel contact, and the $t\to 0$ case models a metal contact. Contact region with area $S_{0}$ is shown by gray color. We assume that $BS_{0}\ll\Phi_{0},$ so that the field does not affect S-matrix.
The simplest scattering matrix of a non-magnetic lead has the form
$\hat{S}^{\mathrm{HI}}=\left(\begin{array}[]{cccc}-t&r&0&0\\ r&t&0&0\\ 0&0&-t&r\\ 0&0&r&t\\ \end{array}\right)\,,\quad t^{2}+r^{2}=1,$ (4)
where two identical blocks are responsible for two spins, and the basis is chosen in accordance with the spin polarization of the helical states at the point of contact (red and blue in Fig. 4). It is convenient to parameterize amplitudes $t$ and $r$ by a parameter $\lambda$:
$t=e^{-\lambda},\quad r=\sqrt{1-e^{-2\lambda}},\quad 0<\lambda<\infty.$ (5)
Due to the topological protection, backscattering appears in the HI only in presence of a defect that violate time-reversal symmetry, in particular, in presence of magnetic defects or in case of magnetization of the leads. We assume in this paper that the leads are non-magnetic, but there is a static backscattering defect in one of the arms of the interferometer. As we already mentioned, much attention was drawn recently to the infinite HES with a dynamic magnetic impurity Väyrynen and Glazman (2017); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022)), whose magnetic moment changes direction after each scattering event. For the isotropic exchange interaction between the impurity and the HES, the impurity’s magnetic moment relaxes within the so-called Korringa time towards the local direction of the electron spin in the HES and, as a consequence, the interaction between the impurity and the HES becomes completely ineffective (see discussion in Väyrynen and Glazman (2017)). Accordingly, the problem of the noise intensity at zero frequency in the HES with a single impurity makes sense only in the presence of an anisotropic exchange interaction Kurilovich et al. (2019a) (or an external magnetic field acting on a dynamic impurity Probst et al. (2022)), while in the limit of isotropic exchange interaction Fano factor is singular Kurilovich et al. (2019a), i.e. the result for the $\mathcal{F}$ depends on the order in which the constants responsible for the anisotropy tend to zero (see also discussion in Nagaev et al. (2018))
At the same time, the relaxation of the impurity’s magnetic moment in reality stems not only from the interaction with the HES, but also from the environment of the impurity, which should provide a non-singular response for the $\mathcal{F}$ even for the isotropic exchange interaction. Therefore, it seems no less interesting to study the case which is opposite to one considered in the papers Väyrynen and Glazman (2017); Kurilovich et al. (2019a, b); Pashinsky et al. (2020); Hsu et al. (2021); Probst et al. (2022), namely, the case of a static magnetic defect with a large spin, which is robustly connected to the external environment.
Such a defect ensures the existence of a magnetic field in a small region of the HES, i.e. it allows elastic backward scattering without tunneling coupling between the HES and defect. Possible experimental realization is, for example, a potential dielectric ferromagnetic point contact with high magnetic stiffness and with the magnetic moment, whose direction is determined by uniaxial anisotropy and the demagnetization tensor of the ferromagnet. The possibility of creating static magnetic contacts to the HES has also been discussed Khomitsky et al. (2022).
Importantly, backscattering can occur not only due to magnetic defect but also due to the interaction with a charged puddle in the bulk of the topological insulator Väyrynen et al. (2013) or due to a point-like non-magnetic scatterer, with account for the electron-electron interaction Sablikov and Sukhanov (2021).
Here, we assume the presence of backscattering static defect (BD) at the edge of the TI, which is described by the scattering matrix of the most general type
$\hat{S}_{\mathrm{BD}}=\left(\begin{array}[]{cc}\cos\theta e^{i\alpha}&i\sin\theta~e^{i\varphi}\\ i\sin\theta~e^{-i\varphi}&\cos\theta e^{-i\alpha}\\ \end{array}\right)\,.$ (6)
For the case of magnetic defect, we neglect the back influence of the HES on the parameters of $S_{\mathrm{BD}}$. The backward scattering rate, $R_{\theta}=\sin^{2}\theta,$ is determined by the quantity $\theta$ while the phase $\varphi$ has the meaning of the backward scattering phase on the BD.
Here we derive analytical expressions for a single BD placed in the upper shoulder. We consider interferometer with the lengths of the upper and lower shoulders given by $L_{1}$ and $L_{2},$ respectively. The BD is placed at position $x_{0}$ such that $0<x_{0}<L_{1}.$ Using expression for scattering matrix (6), one can easily find transfer matrix of the defect
$\hat{T}_{\mathrm{BD}}=\frac{e^{-i\alpha}}{\cos\theta}\left(\begin{array}[]{cc}1&i\sin\theta e^{-i\xi}\\ -i\sin\theta e^{i\xi}&1\end{array}\right),$ (7)
where $\xi=\varphi-2kx_{0}$ and $k=\varepsilon/v_{\rm F}$ is the electron wave vector. One can show that the forward scattering phase $\alpha$ can be fully incorporated into the shift of $\phi,$ so that we put $\alpha=0$ below.
III Shot noise and conductance
III.1 General equations
In this Section, we calculate $\mathcal{F}$ for setups shown in Fig. 1 assuming that a fixed bias voltage $V$ is applied to leads. (In this paper, we assume that the external impedance is zero and therefore the voltage does not fluctuate. For a finite impedance both current and voltage fluctuate as discussed in detail in Ref. Blanter and Büttiker (2000).)
We will consider the most interesting and easily realized case:
$\Delta\ll T\ll eV.$ (8)
The current noise is related with fluctuations of the electric current with respect to its average value $\delta\hat{I}(t)=\hat{I}(t)-\langle\hat{I}\rangle$. Here $\hat{I}$ is the current operator (an analytical expression for $\hat{I}$ is given in de Jong and Beenakker (1997); Blanter and Büttiker (2000)).
The current correlation function associated with the noise is defined by:
${\mathcal{S}}\left(t-t^{\prime}\right)=\frac{1}{2}\langle\delta\hat{I}(t)\delta\hat{I}(t^{\prime})+\delta\hat{I}(t^{\prime})\delta\hat{I}(t)\rangle\,.$
The Fourier transform of $\mathcal{S}$} gives an expression for the noise power: $S(\omega)=2\int_{-\infty}^{\infty}dt\,e^{i\omega t}{S(t)}$ [the factor $2$ in this expression is a matter of convention, see Eq. (1) in Ref. de Jong and Beenakker (1997) and the comment after Eq. (49) in Ref. Blanter and Büttiker (2000)].
Spin-dependent transport through the two-terminal device is fully characterized by the matrix of transmission amplitudes $\hat{t}=t_{\alpha\beta}$ (here $\alpha$ and $\beta$ are the spin indices associated with the outgoing and the incoming electrons, respectively) de Jong and Beenakker (1997); Blanter and Büttiker (2000):
$\mathcal{S}(\omega=0)$ $=2G_{0}$ (9)
$\times\int_{\mu}^{\mu+eV}d\varepsilon\,{\rm Tr}\left[{}\hat{\mathcal{T}}\left(\varepsilon\right)\left(1-\hat{\mathcal{T}}\left(\varepsilon\right)\right)\right]\,$,
where $G_{0}=e^{2}/h$ conductance quantum and
$\hat{\mathcal{T}}(\varepsilon)=\hat{t}(\varepsilon)\,\hat{t}^{\dagger}(\varepsilon).$ (10)
The averaged current, $I=\langle\hat{I}\rangle$, and the Fano factor are given by
$eI=G_{0}\int_{\mu}^{\mu+eV}d\varepsilon\,{\rm Tr}\left[\hat{\mathcal{T}}\left(\varepsilon\right)\right]\,$, (11)
$\mathcal{F}=\frac{\mathcal{S}(\omega=0)}{2eI}=\frac{\int_{\mu}^{\mu+eV}d\varepsilon\,\mathrm{Tr}\left[\hat{\mathcal{T}}(1-\hat{\mathcal{T}})\right]}{\int_{\mu}^{\mu+eV}d\varepsilon\,\mathrm{Tr}[\hat{\mathcal{T}}]}\,$. (12)
The transmission amplitudes $t_{\alpha\beta}(\varepsilon)$ varies on an energy scale on the order of the level spacing. We focus on the case, when the conditions (8) are satisfied. Then for the $\mathcal{F}$ we have
$\mathcal{F}=\frac{\mathrm{Tr}\langle\hat{\mathcal{T}}(1-\hat{\mathcal{T}})\rangle_{\varepsilon}}{\mathrm{Tr}\langle\hat{\mathcal{T}}\rangle_{\varepsilon}}\,$, (13)
where the averaging is taken over a temperature window in the vicinity of the Fermi level in the limit $T\gg\Delta$. In what follows we assume the linearized form of the spectrum, with $\varepsilon=v_{\rm F}k$, so that the energy averaging is reduced to calculation of the integral $\left\langle\cdots\right\rangle_{\varepsilon}=\Delta^{-1}\int_{0}^{\Delta}d\varepsilon\,(\cdots)=\frac{L}{2\pi}\int_{0}^{2\pi/L}dk\,(\cdots)$.
As seen from (11), the conductance is proportional to the transmission coefficient averaged over the spin and the energy:
${\mathcal{T}}=\mathrm{Tr}\langle\hat{\mathcal{T}}\rangle_{\varepsilon}/2.$ (14)
Introducing also the average
${\mathcal{T}}_{2}=\mathrm{Tr}\langle\hat{\mathcal{T}}\hat{\mathcal{T}}\rangle_{\varepsilon},$ (15)
one can write the Fano factor in the following form:
$\mathcal{F}=1-\frac{{\mathcal{T}_{2}}}{2{\mathcal{T}}},$ (16)
In the next two subsections we will discuss calculations that allow one to find $\mathcal{F},$ using Eq. (16). Technically, the key idea is to present the matrix of transmission amplitudes in a special form, which allows performing the energy averaging analytically.
III.2 Conventional AB interferometer
We start with calculation of $\mathcal{F}$ for CI having two arms with the lengths $L/2\pm a$ and two contacts (not necessarily identical) described by scattering matrices with real amplitudes [see Eq.  (2)]. We first find convenient expression for the matrix of transmission amplitude (that consists of one element for spinless CI, also conductance quantum is $G_{0}=e^{2}/h$ in this case), substitute it into the formula (13), and average over the energy.
III.2.1 Method of calculation
We use the method developed in Ref. Shmakov et al. (2013) to sum amplitudes with different winding numbers. We classify all amplitudes of electrons going out of the ring by their direction, either clockwise, $(+)$, or anticlockwise, ()(-). For CI with identical contacts, the amplitudes describing processes of passing the ring in opposite directions without extra revolutions are given by
$\left(\begin{array}[]{c}\beta_{0}^{+}\\ \beta_{0}^{-}\end{array}\right)=t_{\mathrm{in}}t_{\mathrm{out}}\left(\begin{array}[]{c}e^{i(k-2\pi\phi/L)(L/2+a)}\\ e^{i(k+2\pi\phi/L)(L/2-a)}\end{array}\right)\,.$ (17)
All other trajectories finishing at the right lead can be obtained recursively, multiplying by a matrix describing processes which increase number of revolutions, clockwise or counterclockwise, by unity (for a more detailed discussion, see Ref. Shmakov et al. (2013)):
$\left(\begin{array}[]{c}\beta_{n+1}^{+}\\ \beta_{n+1}^{-}\end{array}\right)=\hat{A}\left(\begin{array}[]{c}\beta_{n}^{+}\\ \beta_{n}^{-}\end{array}\right),$ (18)
$\hat{A}=e^{ikL}$
$\times\left(\begin{array}[]{cc}t^{2}e^{-i2\pi\phi}+t_{b}^{2}e^{i2ka}&tt_{b}(e^{-i2\pi\phi}+e^{i2ka})\\ tt_{b}(e^{i2\pi\phi}+e^{-i2ka})&t^{2}e^{i2\pi\phi}+t_{b}^{2}e^{-i2ka}\end{array}\right).$
The full transmission amplitude across the ring is given by:
${t}^{\rm CI}(\varepsilon)=\sum_{n=0}^{\infty}\alpha\hat{A}^{n}\beta_{0}=\alpha(1-\hat{A})^{-1}\beta_{0}\,$, (19)
$\alpha=\left(\begin{array}[]{c}1\\ 1\end{array}\right).$
Having particular expressions for $t^{\rm CI}(\varepsilon),$ we get ${\cal T}^{\mathrm{CI}}(\varepsilon)=|t^{\rm CI}(\varepsilon)|^{2}$ and ${\cal T}_{2}^{\mathrm{CI}}(\varepsilon)=|t^{\rm CI}(\varepsilon)|^{4}$. The averaging over $\varepsilon$ is reduced to integration over the unit circle in complex plane of $z=e^{ikL}$ and is easily performed by residues, see A. Most interesting and physically transparent limiting cases are presented below. In the most general case of interferometer with contacts of different strength, the obtained analytical expressions are rather cumbersome and we present them in the B.
III.2.2 Results (symmetric setup, identical contacts)
Figure 5: The dependence of Fano factor, $\mathcal{F}^{\mathrm{CI}}(\phi,t_{\mathrm{b}})$, on the amplitude of backscattering on the contact, $t_{\mathrm{b}}$, and the magnetic flux, $\phi$, as described by Eqs. (2), (20), (21).
Figure 6: The transmission coefficient (blue curves) and the Fano factor (red curves) for conventional interferometer with two identical contacts. In symmetric interferometer with $a=0$ and weak tunneling coupling ($\gamma=0.1$) there is antiresonance in $\mathcal{T}^{\rm CI}$ and resonance in $\mathcal{F}^{\rm CI}$ at half-integer values of the flux [panels (a) and (b)]. In asymmetric interferometer with $a\neq 0$ and weak tunneling coupling ($\gamma=0.1$, $k_{F}a=1.0$), there also appear peaks at integer values of the flux [panels (c) and (d)]. With increasing $a,$ peaks at $\phi=1/2+n$ decrease in amplitude and become narrower, while peaks at $\phi=n$ grow and become wider (here, n is integer). Shape of the curve $\mathcal{T}^{\rm CI}(\phi)$ strongly depends on the asymmetry of the contacts.
For the case of a symmetric interferometer with equal arm lengths ($a=0$) and identical contacts, the energy-averaged transmission coefficient Dmitriev et al. (2010) is given by
${\mathcal{T}}^{\mathrm{CI}}(\phi,\gamma)=\frac{2\gamma\cos^{2}(\pi\phi)}{\gamma^{2}+\cos^{2}(\pi\phi)}\,.$ (20)
Using formulas from A we find the Fano factor in the form (this equation was presented in Ref. Niyazov et al. (2024) without derivation)
$\mathcal{F}^{\mathrm{CI}}(\phi,\gamma)\!=1\!$ (21)
$-\!\frac{\cos^{2}(\pi\phi)\!\left[5\gamma^{2}\!+\!\gamma^{4}\!+\!(1\!+\!\gamma^{2})\cos^{2}(\pi\phi)\right]}{2\left[\gamma^{2}+\cos^{2}(\pi\phi)\right]^{2}}.$
Equation (21) is illustrated in Fig. 5, where the dependence of Fano factor on the magnetic flux and on the backscattering on the contact, $|t_{\mathrm{b}}|=\gamma/(1+\gamma).$ For three types of tunneling contacts shown in Fig. 3, Eq. (21) simplifies:
$\mathcal{F}^{\mathrm{CI}}(\phi,0)=$ $\frac{1}{2}\,$, (22)
$\mathcal{F}^{\mathrm{CI}}(\phi,\infty)=$ $1-\frac{1}{2}\cos^{2}(\pi\phi)\,$,
$\mathcal{F}^{\mathrm{CI}}(\phi,1)=$ $\frac{4\sin^{2}(\pi\phi)}{(\cos(2\pi\phi)+3)^{2}}\,.$
The value $1/2$ for $\gamma\to 0$ is a standard value for the Fano factor of tunneling level. We also note that the “metallic” contacts with $\gamma=1$, shows both minimal value $\mathcal{F}^{\mathrm{CI}}=0$, at $\phi=0$, and to the maximal value, $\mathcal{F}^{\mathrm{CI}}=1$, at $\phi\to 1/2.$
The dependencies of ${\mathcal{T}}^{\mathrm{CI}}$ and $\mathcal{F}^{\mathrm{CI}}$ on the magnetic flux are shown in Fig. 6a, 6b. The latter figure yields cross-section of Fig. 5 for fixed $\gamma.$ In Fig. 6a we see that in the symmetrical interferometer, $a=0,$ the transmission coefficient and the Fano factor have, respectively, sharp antiresonance and sharp resonance at $\phi=1/2$ (and, hence, at any half-integer flux value).
Figure 7: Parametric plot $(\mathcal{F}^{\rm CI},\mathcal{T}^{\rm CI})$ for fixed $\phi$ and $0<\gamma<\infty$ (a), for fixed $|t_{\rm b}|=\gamma/(1+\gamma)$ and $0<\phi<1$ (b).
Several interesting properties of $\mathcal{F}^{\mathrm{CI}}(\phi,\gamma)$ are to be mentioned here. Firstly,
$\mathcal{F}^{\mathrm{CI}}(0,\gamma)=\mathcal{F}^{\mathrm{CI}}(0,1/\gamma)=\frac{1}{2}\left(\frac{1-\gamma^{2}}{1+\gamma^{2}}\right)^{2},$ (23)
and also
$\mathcal{T}^{\mathrm{CI}}(0,\gamma)=\mathcal{T}^{\mathrm{CI}}(0,1/\gamma)=\frac{2\gamma}{1+\gamma^{2}}.$
It means that for zero flux, the setup with the contacts shown in Fig. 3a with a certain $\gamma$ is indistinguishable in terms of noise and conductance from the setup with the contacts shown in Fig. 3b with $\gamma_{1}=1/\gamma$.
Let us discuss this point in more detail. As seen from Eqs. (20) and (21), at $\phi=0$ the Fano factor and the transmission coefficient are related as follows:
$\mathcal{F}^{\mathrm{CI}}=\frac{1}{2}\left[1-(\mathcal{T}^{\mathrm{CI}})^{2}\right]\,.$ (24)
This equation is shown by red curve in Fig. 7 and is valid both for $\gamma<1$ (i.e. for contacts shown in Fig. 3a) and for $\gamma>1$ (i.e. for contacts shown in Fig. 3b). Specifically, changing $\gamma$ from $\gamma=0$ to $\gamma=\infty,$ we pass the red curve twice: once in the interval $0<\gamma<1$ and once again for $1<\gamma<\infty.$ This is clearly seen from the cross-section of Fig. 5 at $\phi=0.$ Having in mind to compare below with the HI, we note that the role of the amplitude $t_{\mathrm{b}}$ is somewhat similar to the scattering on the magnetic impurity in case of HI, considered below. For the case of CI, dependence on $|t_{\rm b}|$ at $\phi=0$ (i.e. on the red curve Fig. 3a) directly follows from Eq. (23):
$\mathcal{F}^{\rm CI}_{\phi=0}=\frac{1}{2}\frac{(|t_{\rm b}|-1/2)^{2}}{[(|t_{\rm b}|-1/2)^{2}+1/4]^{2}}.$
Importantly, the Fano factor has minimum with zero minimal value in the “metallic” point, $t_{\rm b}=1/2$ ($\gamma=1$). This property is lost in presence of magnetic flux, see Fig. 5.
For $\phi\neq 0,$ the red curve splits into two curves, as shown in Fig.  7 by blue, green and orange lines. Notice that the Fano factor can have two values for the same value of the transmission coefficient, corresponding to two different $\gamma.$ Remarkably, there is a special singular point, $\phi=1/2,$ where the conductance is exactly zero (see Büttiker et al. (1984) and a more detailed discussion of this property in Dmitriev et al. (2010)). At this point, the Fano factor becomes unity for fixed $\gamma$:
$\mathcal{T}^{\mathrm{CI}}(1/2,\gamma)=0,\quad\mathcal{F}^{\mathrm{CI}}(1/2,\gamma)=1.$ (25)
The behavior of the Fano factor near this point is illustrated by the orange curve in Fig. 7a, which corresponds to $\phi=0.48.$ The transmission coefficient and the Fano factor for $|\delta\phi|=|\phi-1/2|\ll 1,$ weak tunneling coupling, $\gamma\ll 1,$ and arbitrary ratio
$x={\pi\delta\phi}/{\gamma}.$
are given by the following expressions
$\mathcal{T}^{\rm CI}(\delta\phi,\gamma)$ $\approx 2\gamma\frac{x^{2}}{1+x^{2}},$ (26)
$\mathcal{F}^{\mathrm{CI}}(\delta\phi,\gamma)$ $\approx\frac{2-x^{2}+x^{4}}{2(1+x^{2})^{2}}\,$, (27)
Hence, the value of $\mathcal{F}^{\mathrm{CI}}$ depends on the order of taking the limits $\delta\phi\to 0$ and $\gamma\to 0,$ so that if we fix $\delta\phi$ and tend $\gamma$ to zero, then instead of Eq. (25) we get the usual expression for the tunnel contact:
$\mathcal{F}^{\mathrm{CI}}\to 1/2,\quad\gamma\to 0,\quad\delta\phi~\mbox{ is fixed}.$
From Eqs. (27) we can get an analytical expression describing the orange line in Fig. 7a. To this end, we fix $\delta\phi$ and change tunneling coupling from $\gamma=0$ to relatively large value: $\delta\phi\ll\gamma\ll 1.$ Then, we get at $0<\mathcal{T}^{\mathrm{CI}}<\pi|\delta\phi|$:
$\mathcal{F}^{\mathrm{CI}}\approx\frac{3}{4}-\frac{1}{2}\frac{\mathcal{T}^{\mathrm{CI}}}{\pi\delta\phi}\pm\frac{1}{4}\sqrt{1-\left(\frac{\mathcal{T}^{\mathrm{CI}}}{\pi\delta\phi}\right)^{2}}\,.$ (28)
The sign in front of the last term distinguish two branches of the orange curve. The minimum of $\mathcal{T}^{\rm CI}$ is reached at the lower branch at $\mathcal{T}^{\rm CI}=15\pi|\delta\phi|/16$ and is given by $\mathcal{F}^{\rm CI}=17/32,$ , which is the value of the dashed line in Fig. 7a for $\mathcal{T}^{\mathrm{CI}}=0$.
In Fig. 7a we fixed $\phi$ and changed $\gamma$ from zero to infinity. Fixing $\gamma$ instead and changing $\phi$ in the interval $0<\phi<1$, we find the following expression connecting transmission coefficient and the Fano factor valid for any $\gamma$ in the interval $0<\gamma<\infty$:
$\mathcal{F}^{\mathrm{CI}}$ $=$ $1-\frac{5+\gamma^{2}}{2}~\frac{\mathcal{T}^{\mathrm{CI}}}{2\gamma}+2\left(\frac{\mathcal{T}^{\mathrm{CI}}}{2\gamma}\right)^{2},$
$0<\mathcal{T}^{\mathrm{CI}}<\frac{2\gamma}{1+\gamma^{2}}.$
This equation has minimum at $\mathcal{T}^{\mathrm{CI}}_{\rm min}=(5+\gamma^{2})\gamma/4$ given by $\mathcal{F}^{\mathrm{CI}}_{\rm min}=[7-\gamma^{2}(10+\gamma^{2})]/32.$ The family of the curves described by Eq. (III.2.2) for different $\gamma$ is shown in Fig. 7b. Note that the dashed curves in panels (a) and (b) of Fig. 7 coincide. From above discussion, we see that behavior of the noise of CI is very singular for $\phi$ close to $1/2.$ This means that noise is very sensitive to the flux in the vicinity of this point.
III.2.3 Results (asymmetric setup)
The asymmetry of setup becomes important when $a\geq k_{F}^{-1}.$ For not too large voltages, $eV\ll v_{F}/a$, one can replace $ka\rightarrow k_{F}a$ in scattering amplitudes. The expression for the transmission coefficient in asymmetric setup was obtained earlier in Ref. Dmitriev et al. (2010). Analytical expression for the Fano factor in this case can be found by using formulas presented in A and B. Its general form is too cumbersome and we present it here only for the flux values close to integer and half-integer number and for weak tunneling coupling.
For $\delta\phi=|\phi-1/2|\ll 1$ and $\gamma\ll 1$ the Fano factor reads:
$\mathcal{F}^{\mathrm{CI}}_{1/2}(\delta\phi,\gamma)$ $=\frac{2-\bar{x}^{2}+\bar{x}^{4}}{2(1+\bar{x}^{2})^{2}}$ (30)
$+\frac{\sin^{2}(k_{F}a)[3\bar{x}^{2}-\sin^{2}(k_{F}a)]}{2(1+\bar{x}^{2})[\bar{x}^{2}+\sin^{2}(k_{F}a)]},$
where
$\bar{x}=\frac{\pi\,\delta\phi}{\gamma\cos(k_{F}a)}.$
Similar to Eq. (27), the value of $\mathcal{F}^{\mathrm{CI}}_{1/2}$ at $\delta\phi\to 0$, $\gamma\to 0$ depends on the order of limits. Fixing $\delta\phi$ and decreasing the tunneling coupling to zero, $\gamma\to 0,$ we get $\bar{x}\to\infty,$ so that $\mathcal{F}^{\mathrm{CI}}_{1/2}\to 1/2$. Putting $\delta\phi$ to zero first, we find
$\mathcal{F}^{\mathrm{CI}}_{1/2}\to 1-\frac{\sin^{2}(k_{\rm F}a)}{2}.$ (31)
It turns out, that the Eq. (30) also describes the case of small magnetic flux, $|\phi|\ll 1$, if we set $\delta\phi=\phi$ there and replace $k_{F}a$ by $k_{F}a+\pi/2$.
As seen from Eqs. (30) and (31), the asymmetry of the device strongly affect the noise intensity even for fixed magnetic flux, so that changing $a$ within a small interval about the Fermi wavelength, we change $\mathcal{F}^{\rm CI}$ by a factor on the order of two.
The dependencies of ${\mathcal{T}}^{\mathrm{CI}}$ and $\mathcal{F}^{\mathrm{CI}}$ on the magnetic flux, $\phi$, are shown in Fig. 6c and 6d. Comparing these figures with Fig. 6a and 6b plotted for symmetric case, we see that new peaks at integer values of the flux appear for $a\neq 0$ both in the transmission coefficient (antiresonances) and in the Fano factor (resonances) for integer values of flux. With increasing asymmetry, the amplitudes of the peaks corresponding to half-integer values of the flux decrease, while the amplitudes of the peaks at integer flux values increase. We conclude that asymmetry of the device, which is always present in real experiments Shea et al. (2000); Piazza et al. (2000); Fuhrer et al. (2001); Keyser et al. (2003); Zou et al. (2007), strongly modifies shot noise.
III.2.4 Results (different tunneling couplings)
In this subsection, we present results for the case of symmetric setup, $a=0,$ but different right and left contacts, characterized by parameter, $\gamma_{\mathrm{R}}$ and $\gamma_{\mathrm{L}}$, respectively. Using general equations presented in B and performing standard energy averaging with the use of formulas from A, one can find the following expressions for the transmission coefficient and Fano factor:
${\mathcal{T}}^{\mathrm{CI}}(\phi,\gamma_{\mathrm{R}},\gamma_{\mathrm{L}})=\kappa^{-1/2}\frac{2\bar{\gamma}\cos^{2}(\pi\phi)}{\bar{\gamma}^{2}+\cos^{2}(\pi\phi)}\,$,
$\mathcal{F}^{\mathrm{CI}}(\phi,\gamma_{\mathrm{R}},\gamma_{\mathrm{L}})=1-\kappa^{-1}\frac{\cos^{2}(\pi\phi)}{2(\bar{\gamma}^{2}+\cos^{2}(\pi\phi))^{2}}$
$\times\left((1+4\kappa)\bar{\gamma}^{2}+\bar{\gamma}^{4}+(1+\bar{\gamma}^{2})\cos^{2}(\pi\phi)\right)\,$,
where
$\bar{\gamma}=\sqrt{\gamma_{\mathrm{L}}\gamma_{\mathrm{R}}}\,,\quad\kappa=(\gamma_{\mathrm{L}}+\gamma_{\mathrm{R}})^{2}/4\gamma_{\mathrm{L}}\gamma_{\mathrm{R}}\,.$
The analog of Eq. (24) at $\phi=0$ for different right and left contacts attains the form
$\mathcal{F}^{\mathrm{CI}}=1-\frac{1}{2\kappa}-\frac{\kappa}{2}\left({\mathcal{T}}^{\mathrm{CI}}\right)^{2}\,.$ (32)
III.3 Helical AB interferometer
Next, we discuss the Fano factor of the current noise in HI having two arms with lengths $L_{1}$ and $L_{2}.$ Just as in the case of CI we start formula for energy dependent amplitude $\hat{t}^{\rm HI}(\varepsilon),$ which is a matrix in spinful case, and then perform energy averaging. The analytical expression obtained for $\mathcal{F}^{\rm HI}$ is rather cumbersome, so we present it in the C focusing in the main text on the most interesting limiting cases.
III.3.1 Scattering matrix
The matrix of the transmission amplitudes $\hat{t}^{\rm HI}$ from one contact to another is defined as follows
$\left(\begin{array}[]{c}a^{\uparrow}\\ a^{\downarrow}\end{array}\right)=\hat{t}^{\rm HI}\left(\begin{array}[]{c}b^{\uparrow}\\ b^{\downarrow}\end{array}\right)\,$, (33)
where $(b^{\uparrow},b^{\downarrow})$ and $(a^{\uparrow},a^{\downarrow})$ are the amplitudes of incoming (from the right contact) and outgoing (to the left contact) waves, respectively (see Fig. 1). This matrix has been obtained earlier Niyazov et al. (2020):
$\hat{t}^{\rm HI}$ $=\frac{r^{2}e^{2\pi i\phi L_{1}/L}}{t^{2}}\left(\begin{array}[]{cc}e^{ikL_{1}}&0\\ 0&e^{-ikL_{1}}\end{array}\right)$ (34)
$\times\left(\begin{array}[]{cc}t&0\\ 0&1\end{array}\right)\hat{g}\left(\begin{array}[]{cc}t&0\\ 0&1\end{array}\right)\,$,
$\hat{g}$ $=\cos\theta\left[\left(\begin{array}[]{cc}0&0\\ 0&-1\end{array}\right)\right$.
$\left.+\frac{1}{2}\sum\limits_{\alpha=\pm 1}\frac{1+\alpha\hat{H}}{1-t^{2}e^{i(kL+\alpha 2\pi\phi_{0})}}\right]\,.$
Here, $t$ and $r$ are scattering amplitudes entering matrix Eq. (4), $\phi_{0}$ is determined by the relation
$\cos(2\pi\phi_{0})=\cos\theta\cos(2\pi\phi)\,$, (35)
and the matrix $\hat{H}$ is given by
$\hat{H}=\left(\begin{array}[]{cc}a&be^{i\xi}\\ be^{-i\xi}&-a\end{array}\right)\,$, (36)
where $\xi$ is defined in Eq. (7). The coefficients
$a=i\frac{e^{-2\pi i\phi}-\cos(2\pi\phi_{0})\cos\theta}{\cos\theta\sin(2\pi\phi_{0})}\,$, (37)
$b=\frac{e^{-2\pi i\phi}\tan\theta}{\sin(2\pi\phi_{0})}\,$, (38)
are related by the relation $a^{2}+b^{2}=1.$
Calculations shows that asymmetry of the arms does not play any role in the HI. Also, position of MD $x_{0}$ drops out from final expressions for $\mathcal{T}^{\rm HI}$ and $\mathcal{F}^{\rm HI}.$ We therefore consider two cases: identical contacts (see next subsection and C.1) and non-identical contacts (see C.2)
Generalization of Eq. (34) for the case of different tunneling contacts characterized by amplitudes $t_{\rm L}$ and $t_{\rm R}$ is straightforward. One should make in this equation the following replacement:
$t\to\sqrt{t_{\rm L}t_{\rm R}},\quad r\to\sqrt{r_{\rm L}r_{\rm R}},\quad\xi\to\xi+i\beta,$
where the parameter describing the difference of tunneling contacts [see also C.2] is
$\beta=\frac{1}{2}\ln(t_{\mathrm{R}}/t_{\mathrm{L}})\,.$ (39)
Figure 8: Panel (a): The dependence of the Fano factor, $\mathcal{F}^{\mathrm{HI}}$, on $r^{2}$ and the strength of the magnetic impurity, $R_{\theta}$; for $r\to 1$ there an is optimal value, $R_{\theta}=R_{\theta}^{\rm opt},$ where the Fano factor has maximum. Panel (b): Cross-sections of panel (a) at different values of $R_{\theta};$ dependence on the flux is weak and is illustrated for $R_{\theta}=R_{\theta}^{\rm opt}$ by broadening of the curve (shown by blue color) for flux belonging to interval $0<\phi<1$. (these figures were presented in Ref. Niyazov et al. (2024) in a slightly different way). The point $r=1$ is “noiseless” both for ballistic case $R_{\theta}=0$ and the case of very strong defect $R_{\theta}=1.$
III.3.2 Results (identical contacts)
The transmission coefficient and the Fano factor are obtained with the use of Eqs. (10), (14), (15), (16), and (34). Direct energy averaging in Eqs. (14) and (15) yields Niyazov et al. (2021a)
${\mathcal{T}}^{\mathrm{HI}}$ $=\tanh\lambda\left[1\right$. (40)
$\left.-\frac{\sin^{2}\theta\sinh^{2}\lambda\cosh(2\lambda)}{\cosh^{2}(2\lambda)-\cos^{2}\theta\cos^{2}(2\pi\phi)}\right]$
and Niyazov et al. (2024)
${\mathcal{T}}_{2}^{\mathrm{HI}}$ $=\tilde{C}\sum_{m,n=0}^{3}\tilde{A}_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,.$ (41)
Here
$D=$ $\sinh^{2}(2\lambda)+R_{\theta}+(1-R_{\theta})R_{\phi}\,$, (42)
$\tilde{C}$= $\tanh\lambda/\cosh^{2}\lambda,$
$R_{\theta}=$ $\sin^{2}\theta,\quad R_{\phi}=\sin^{2}(2\pi\phi),$
and the coefficients $\tilde{A}_{(m,n)}$ depend on $R_{\theta},R_{\phi}$ do not depend on $\lambda.$ Analytical expressions for these coefficients are presented in the C.1. The Fano factor is connected with $\mathcal{T}^{\rm HI}$ and $\mathcal{T}^{\rm HI}_{2}$ by Eq. (16).
In Fig. 8a, we present dependence of the Fano factor, on $r^{2}$ and the strength of the magnetic impurity, $R_{\theta}$ for $\phi=0$. For “metallic” contact, $r\to 1,$ dependence on the backscattering strength $R_{\theta}$ is non-monotonous. Comparing this figure with analogous Fig. 5 for CI, we see that in both cases the Fano factor has extremum as a function of squared backscattering amplitude. However, for CI the Fano factor has minimum, while for HI it has maximum.
In Fig. 8b, we plot dependence of $\mathcal{F}^{\rm HI}$ on $r^{2}$ for fixed $R_{\theta}.$ Two limiting dependencies, corresponding to absence of impurity, $R_{\theta}=0,$ and very strong impurity, $R_{\theta}=1,$ are shown by orange and green line, respectively. In both cases, there is no dependence on the magnetic flux. For any other $R_{\theta}$ in the interval $0<R_{\theta}<1,$ the Fano factor depends on $\phi$ but this dependence is much weaker as compared to CI. Each curve corresponding to certain $R_{\theta}$ broadens as illustrated by blue curve corresponding to $R_{\theta}=R_{\theta}^{\rm opt}.$
In order to emphasize this dependence one can consider normalized value
$\mathcal{F}_{\rm n}^{\mathrm{HI}}(\phi)=\frac{\mathcal{F}^{\mathrm{HI}}(\phi)-\mathcal{F}^{\mathrm{HI}}(1/4)}{\mathcal{F}^{\mathrm{HI}}(0)-\mathcal{F}^{\mathrm{HI}}(1/4)}.$ (43)
The dependence of the normalized quantity $\mathcal{F}_{\rm n}^{\mathrm{HI}}$ on the magnetic flux for several sets of parameters $R_{\theta}$ and $\lambda$ is shown in Fig. 9 along with the flux dependence of the normalized transmission coefficient. One can see sharp resonances both in $\mathcal{F}_{\rm n}^{\mathrm{HI}}$ and $\mathcal{T}_{\rm n}^{\mathrm{HI}}$ at weak tunneling coupling and small $R_{\theta}$ (see Fig. 9c) which evolve into harmonic Aharonov-Bohm oscillations with increasing $\lambda$ and $R_{\theta}$ (see Fig. 9d)
Figure 9: Dependence of the normalized factor Fano, $\mathcal{F}_{n}^{\mathrm{HI}}$, and the normalized conductance, $\mathcal{T}_{n}^{\mathrm{HI}}$, for the helical interferometer on magnetic flux, $\phi$, at different strengths of scattering by a magnetic defect $R_{\theta}$ and tunneling amplitude, $r$. (a) $r=0.7,R_{\theta}=1/8$ (b) $r=0.7,R_{\theta}=3/8$ (c) $r=0.4,R_{\theta}=1/1000$ (d) $r=0.7,R_{\theta}=1/2$. Shape of the curve $\mathcal{F}^{\rm HI}(\phi)$ strongly depends on the transparency of the contacts.
One can also plot dependence of the Fano factor on the transmission coefficient similar to Fig. 7 for CI. Analogous dependence for HI is shown in Fig. 10. Interesting difference of two pictures is that CI has a single noiseless point, $\mathcal{F}^{\rm CI}=0,$ corresponding to ideal transmission, $\mathcal{T}^{\rm CI}=1,$ while HI in addition to the analogous point with $\mathcal{F}^{\rm HI}=0,~\mathcal{T}^{\rm HI}=1$ has one more noiseless point, corresponding to $R_{\theta}=1,~\lambda=\infty.$ At this point $\mathcal{F}^{\rm HI}=0,~\mathcal{T}^{\rm HI}=1/2.$
Fig. 10 is one of the key results concerning HI. Indeed, measuring simultaneously both conductance and noise and using Fig. 10), one can find the strength of the backscattering defect. It is worth stressing that measuring the conductance only is insufficient to find the strength of the defect, because the conductance depends also on the properties of the contacts, which are actually not known to a sufficient precision. Importantly, by using both conductance and noise measurements, together with Fig. 10 we can find $R_{\theta}$ without having detailed information about amplitudes $t$ and $r,$ characterizing the contacts. It can help to identify backscattering mechanism that apparently always exists in the realistic experimental structures Olshanetsky et al. (2023).
III.3.3 Limiting cases
Here, we discuss several limiting cases allowing for simple analytical description.
In the absence of a magnetic defect, $R_{\theta}=0$, interference effects are absent as we discussed in the introduction, so that expressions for the Fano factor and transmission coefficient dramatically simplify:
$\mathcal{T}^{\rm HI}=\tanh[\lambda],\quad\mathcal{F}^{\rm HI}=\frac{1}{2\cosh^{2}[\lambda]}.$ (44)
These equations not depend on $\phi$ and are related by the same equation as for ballistic CI [compare with Eq. (24)]
$\mathcal{F}^{\mathrm{HI}}=\frac{1}{2}\left[{1-({\mathcal{T}}^{\mathrm{HI}})^{2}}\right].$ (45)
This dependence is shown in Fig. 10 by red line
The formulas for the conductance and Fano factor in case of different tunneling contacts can also be obtained in closed form and are presented in C.2. An analog of Eq. (45) in case of different tunneling contacts, in contrast to CI case, cf. Eq. (32) looks a bit more cumbersome
$\mathcal{F}^{\rm HI}=\left[1-({\mathcal{T}}^{\rm HI})^{2}\right]\sqrt{\sinh^{2}(2\beta)+({\mathcal{T}}^{\rm HI})^{2}}$ (46)
$/\left(\sqrt{\sinh^{2}(2\beta)+({\mathcal{T}}^{\rm HI})^{2}}+{\mathcal{T}}^{\rm HI}\cosh(2\beta)\right)\,$,
with the asymmetry parameter $\beta=\frac{1}{2}\ln(t_{\mathrm{R}}/t_{\mathrm{L}})$, Eq. (39).
Figure 10: The combined dependence of Fano factor, $\mathcal{F}^{\mathrm{HI}}$, and the transmission coefficient, $\mathcal{T}^{\mathrm{HI}}$, on the transparency of the contact, $\lambda$, for different strength of the magnetic impurity, $R_{\theta}$, and the magnetic flux $\phi=0.4$, according to Eqs. (40), (41). The black dashed line borders the area of maximum transmission coefficient for given $R_{\theta}$.
The Fano factor and transmission coefficient also do not depend on the magnetic flux for the strong magnetic impurity, $R_{\theta}=1$ (green curves on the Fig. 8, accordingly). In this case, we have
$\mathcal{T}^{\rm HI}=\frac{1}{2}{\tanh[2\lambda]},\quad\mathcal{F}^{\rm HI}=\frac{1}{2\cosh^{2}[2\lambda]},$ (47)
$\mathcal{F}^{\rm HI}=\frac{1}{2}-2\left(\mathcal{T}^{\rm HI}\right)^{2}.$ (48)
Last dependence is shown by green line in Fig. 10.
The flux dependence is also absent for the open interferometer, $r=1~(\lambda=\infty)$. In this case, we get
$\mathcal{T}^{\mathrm{HI}}=1-\frac{R_{\theta}}{2},$ (49)
$\mathcal{F}^{\mathrm{HI}}=R_{\theta}\frac{1-R_{\theta}}{2-R_{\theta}}=\frac{(2\mathcal{T}^{\mathrm{HI}}-1)(1-\mathcal{T}^{\mathrm{HI}})}{\mathcal{T}^{\mathrm{HI}}}\,.$
Hence, the Fano factor has a maximum as a function of backscattering probability as is shown by the black curve in the Fig. 8 (b).
Equation (49) has a simple physical meaning. For $r=1,$ the windings are absent, and there are two parallel conductors, i.e. lower and upper arms of the HI. The lower arm, with spin $\downarrow$, is an ideal conductor with the transmission coefficient $\mathcal{T}_{\downarrow}=1$ while the transmission through the upper arm is controlled by the magnetic defect, $\mathcal{T}_{\uparrow}=\cos^{2}\theta,$ varying from 0 to $1.$ Then, we get
$\mathcal{F}^{\rm HI}=\frac{\mathcal{T}_{\uparrow}(1-\mathcal{T}_{\uparrow})}{1+\mathcal{T}_{\uparrow}}.$ (50)
Taking into account that the total transmission coefficient in this case is given by $\mathcal{T}^{\mathrm{HI}}=(\mathcal{T}_{\uparrow}+1)/2,$ we arrive to Eq. (49). We have maximum $\mathcal{F}_{\mathrm{max}}^{\rm HI}=3-2\sqrt{2}\approx 0.17$ at $\mathcal{T}_{\uparrow,\mathrm{max}}=\sqrt{2}-1$ and, respectively, $R_{\theta}^{\mathrm{max}}=2-\sqrt{2}\approx 0.59$. The total transmission coefficient at this point is given by $\mathcal{T}^{\mathrm{HI}}_{\rm max}=1/\sqrt{2}.$ We notice that Eq. (49) describes both black line in Fig. 8a and dashed black line in Fig. 10.
In order, to combine here analytical formulas for all relevant cases, we also present at the end of this section several equations obtained previously in Ref. Niyazov et al. (2024). For almost metallic contact, $t\ll 1$ ($\lambda\to\infty$), and for an arbitrary scattering strength on a magnetic defect, $\mathcal{F}^{\rm HI}$ shows weak oscillations with $\phi$:
$\mathcal{F}^{\rm HI}\approx\frac{R_{\theta}(1-R_{\theta})}{2-R_{\theta}}+2t^{2}$ (51)
$+2t^{4}\frac{R_{\theta}(1-R_{\theta})\left(10-12R_{\theta}+3R_{\theta}^{2}\right)}{(2-R_{\theta})^{2}}\cos(4\pi\phi).$
More interesting is the opposite case of an almost tunnel contact: $t\to 1,r=\sqrt{1-t^{2}}\ll\theta.$ In this case,
$\mathcal{F}^{\rm HI}\approx\frac{1}{2}-\frac{r^{4}}{8}\left[1+\frac{3R_{\theta}}{R_{\phi}+R_{\theta}-R_{\theta}R_{\phi}}\right].$ (52)
This formula shows weak but sharp resonances with magnetic flux. These resonances are well pronounced in the normalized Fano factor. Indeed, from Eqs.  (43) and (52) for $\theta\ll 1$ we obtain
$\mathcal{F}_{\rm n}^{\mathrm{HI}}\approx\frac{\theta^{2}}{\theta^{2}+\sin^{2}(2\pi\phi)},$
i.e. sharp resonances at $\phi=0$ and $\phi=1/2.$
IV Conclusion
In the previous sections we analyzed two types of quantum interferometers based on conventional materials and helical edge states. We have demonstrated that the shot noise in both systems is substantially modified by quantum interference effects. It is essential that the interference effects survive in the high temperature limit, $T\gg\Delta.$ The appearance of temperature-robust interference effects stems from trajectories interfering at any energy. In case of symmetric CI for any trajectory entering the interferometer and moving clockwise there is another trajectory moving counterclockwise, and these trajectories interfere destructively at $\phi=1/2.$ Such destructive interference is exact, i.e. the transmission amplitude is identically zero at all energies, $t(\epsilon,\phi=1/2)\equiv 0$ Büttiker et al. (1984) (see also the discussion of the consequences of this identity in Dmitriev et al. (2010, 2015)). The peculiarity of interference effects in HI lies in the fact that they appear only in presence of backscattering defects. In this case trajectories interfering at any energy appear as well, namely, trajectories returning to the defect after $n$ revolutions clockwise and counterclockwise are always interfering. The relative phase of such trajectories is $4\pi n\phi$, and as a result the Aharonov-Bohm oscillation period in HI equals to $\Delta\phi=1/2,$ in contrast to CI, where the oscillation period is equal to $\Delta\phi=1.$
To conclude, in this paper we discuss and compare noise characteristics of two types of Aharonov-Bohm interferometers, based on conventional materials and on helical edges states of TI. We focus on the case of relatively high temperatures, which partially smear the interference effects, all the while leaving many remarkable quantum features intact. In particular, we show that even in case of higher temperatures the noise Fano factor reveals a characteristic periodicity in the magnetic field. We demonstrate different patterns of this periodicity with resonances and anti-resonances at integer and half-integer values of the flux (see Fig. 6 for CI and Fig. 9 for HI).
One of the most important conclusions following from our calculations is that the interference effects are much more pronounced for CI and, consequently, the influence of the magnetic field on the Fano factor is much stronger as compared to HI. The reason is that the interference comes into play for CI already in the ballistic case, when impurities in the arms of CI are absent, while the Aharonov-Bohm peaks in HI are proportional to the backscattering probability.
We show that asymmetrical placements of contacts results in strong modification of noise Fano factor in CI in contrast to HI. In particular, as seen from Eqs. (30) and (31), the asymmetry of the CI strongly affects the noise intensity even for fixed magnetic flux, so that changing difference of shoulder’s lengths, $a,$ within a small interval about the Fermi wavelength, we change $\mathcal{F}_{\rm CI}$ by a factor on the order of unity. We also find that the characteristics of the contact between the CI and the leads strongly modify the noise in this case, even in the ballistic regime inside the CI. In turn, the noise in case of HI is strongly impacted by the backscattering defects inside the HI ring.
We also demonstrate that the simultaneous measurement of conductance and Fano factor in HI allows one to determine the strength of backscattering defects that violate topological protection in the realistic setups. As for CI, it can show different noise for the same transmission coefficient (see, for example, Fig. 7). Hence, measuring, complementary to conductance, also the Fano factor of CI would help to experimentally determine the type of the contacts to leads (see Fig. 3a and Fig. 3b) and the strength of tunneling coupling.
Data availability statement
All data that support the findings of this study are included within the article (and any supplementary files).
Acknowledgments
The work was carried out with financial support from the Russian Science Foundation (grant No. 25-12-00212), https://rscf.ru/en/project/25-12-00212/.
Appendix A Energy averaging
Throughout the paper we average observable quantities over a small temperature window in the vicinity of the Fermi level in the limit $T\gg\Delta$. For the linearized form of the spectrum, with $\varepsilon=v_{F}k$, the energy averaging is reduced at high temperatures to calculating the integral $\left\langle\cdots\right\rangle_{\varepsilon}=\Delta^{-1}\int_{0}^{\Delta}d\varepsilon\,(\cdots)=\frac{L}{2\pi}\int_{0}^{2\pi/L}dk\,(\cdots)$. The averaging over $kL$, is reduced to integration over the unit circle $z=e^{ikL}$ in complex plane and is easily performed by residues. The formulas which were used for calculations are listed below:
[…]
$\left\langle\frac{1}{(1-\tau_{1}e^{\alpha ikL})}\frac{1}{(1-\tau_{2}e^{-\alpha ikL})}\right\rangle_{\varepsilon}$ $=\frac{1}{1-\tau_{1}\tau_{2}}\,$,
$\left\langle\frac{1}{(1-\tau_{1}e^{\alpha ikL})}\frac{1}{(1-\tau_{2}e^{-\alpha ikL})}\frac{1}{(1-\tau_{3}e^{-\alpha ikL})}\right\rangle_{\varepsilon}$ $=\frac{1}{(1-\tau_{1}\tau_{2})(1-\tau_{1}\tau_{3})}\,$,
$\left\langle\frac{1}{(1-\tau_{1}e^{\alpha ikL})}\frac{1}{(1-\tau_{2}e^{-\alpha ikL})}\frac{1}{(1-\tau_{3}e^{-\alpha ikL})}\frac{1}{(1-\tau_{4}e^{-\alpha ikL})}\right\rangle_{\varepsilon}$ $=\frac{1}{(1-\tau_{1}\tau_{2})(1-\tau_{1}\tau_{3})(1-\tau_{1}\tau_{4})}\,$,
$\left\langle\frac{1}{(1-\tau_{1}e^{\alpha ikL})}\frac{1}{(1-\tau_{2}e^{\alpha ikL})}\frac{1}{(1-\tau_{3}e^{-\alpha ikL})}\frac{1}{(1-\tau_{4}e^{-\alpha ikL})}\right\rangle_{\varepsilon}$
$=\frac{1-\tau_{1}\tau_{2}\tau_{3}\tau_{4}}{(1-\tau_{1}\tau_{3})(1-\tau_{2}\tau_{3})(1-\tau_{1}\tau_{4})(1-\tau_{2}\tau_{4})}\,$,
where $\alpha=\pm 1$ and all $|\tau_{j}|<1$.
Appendix B Conventional interferometer with non-equal arms and non-identical contacts
In this Appendix, we present generalization of the tunneling amplitude $t^{\rm CI}(\varepsilon)$ [see Eq. (19)] for the case of non-identical contacts having different tunneling couplings. The symmetric case is obtained by putting $t_{\mathrm{in,L}}=t_{\mathrm{in,R}}=t_{\mathrm{in}}$, etc. in equations presented below.
$t^{\mathrm{CI}}(\varepsilon)=$ $\sum_{n=0}^{\infty}\vec{\alpha}\hat{A}^{n}\vec{\beta}_{0}\,,\quad\vec{\alpha}=\left(\begin{array}[]{c}1\\ 1\end{array}\right)\,,\quad\vec{\beta}=\left(\begin{array}[]{c}\beta_{0}^{+}\\ \beta_{0}^{-}\end{array}\right)=t_{\mathrm{in,L}}t_{\mathrm{out,R}}\left(\begin{array}[]{c}e^{i(k-2\pi\phi/L)(L/2+a)}\\ e^{i(k+2\pi\phi/L)(L/2-a)}\end{array}\right)\,$, (54)
$\hat{A}$= $e^{ikL}\left(\begin{array}[]{cc}t_{\mathrm{L}}t_{\mathrm{R}}e^{-i2\pi\phi}+t_{\mathrm{b,L}}t_{\mathrm{b,R}}e^{i2ka}&t_{\mathrm{L}}t_{\mathrm{b,R}}e^{-i2\pi\phi}+t_{\mathrm{b,L}}t_{\mathrm{R}}e^{i2ka}\\ t_{\mathrm{L}}t_{\mathrm{b,R}}e^{i2\pi\phi}+t_{\mathrm{b,L}}t_{\mathrm{R}}e^{-i2ka}&t_{\mathrm{L}}t_{\mathrm{R}}e^{i2\pi\phi}+t_{\mathrm{b,L}}t_{\mathrm{b,R}}e^{-i2ka}\end{array}\right)\,.$
After some algebra one can rewrite expression for $t^{\rm CI}(\varepsilon)$ as follows:
$t(\varepsilon)^{\rm CI}=$ $t_{\mathrm{in,L}}t_{\mathrm{out,R}}e^{-i2\pi\phi a/L}\left(\frac{Ze^{i(kL/2-\pi\phi)}}{1-\tilde{t}^{2}e^{i{kL-2\pi\phi}}}+\frac{Z^{\ast}|_{\mathrm{L}\leftrightarrow\mathrm{R}}e^{i(kL/2+\pi\phi)}}{1-\tilde{t}^{\ast 2}e^{i{kL+2\pi\phi}}}\right)\,$, (55)
$Z=$ $i\frac{Z_{1}Z_{2}}{Z_{3}}e^{i\pi\phi}\,$,
$Z_{1}=$ $Z_{4}-(t_{\mathrm{b,L}}-t_{\mathrm{L}})(t_{\mathrm{b,R}}\sin(2ak)+t_{\mathrm{R}}\sin(2\pi\phi))-i(t_{\mathrm{b,R}}t_{\mathrm{L}}\cos(2ak)+t_{\mathrm{b,L}}t_{\mathrm{R}}\cos(2\pi\phi))\,$,
$Z_{2}=$ $Z_{4}+(t_{\mathrm{b,R}}-t_{\mathrm{R}})(t_{\mathrm{b,L}}\sin(2ak)+t_{\mathrm{L}}\sin(2\pi\phi))-i(t_{\mathrm{b,L}}t_{\mathrm{R}}\cos(2ak)+t_{\mathrm{b,R}}t_{\mathrm{L}}\cos(2\pi\phi))\,$,
$Z_{3}=$ $2Z_{4}\left(t_{\mathrm{b,L}}t_{\mathrm{R}}e^{-i(ak+\pi\phi)}+t_{\mathrm{b,R}}t_{\mathrm{L}}e^{i(ak+\pi\phi)}\right)\,$,
$Z_{4}^{2}=$ $\left(t_{\mathrm{b,L}}^{2}-t_{\mathrm{L}}^{2}\right)\left(t_{\mathrm{b,R}}^{2}-t_{\mathrm{R}}^{2}\right)-(t_{\mathrm{b,L}}t_{\mathrm{b,R}}\cos(2ak)+t_{\mathrm{L}}t_{\mathrm{R}}\cos(2\pi\phi))^{2}\,$,
$\tilde{t}^{2}=$ $iZ_{4}+t_{\mathrm{b,L}}t_{\mathrm{b,R}}\cos(2ak)+t_{\mathrm{L}}t_{\mathrm{R}}\cos(2\pi\phi)\,.$
Assuming that $eV\ll v_{F}/a$, one can replace $ka\rightarrow k_{F}a$ in this equation. Now we see that the expressions $|t^{\rm CI}(\varepsilon)|^{2}$, $|t^{\rm CI}(\varepsilon)|^{4}$ contain the complex quantity $e^{ikL}$ which is regarded as a new variable. Then, one can perform averaging of these expressions over energy by using formulas derived in A.
Appendix C Helical case
C.1 Identical contacts
General formulas for the conductance and the noise intensity are obtained by averaging the expressions  (9), (10), (11) and (12) in the main text over the energy, using expressions for the amplitude (34). This result was obtained early in Ref. Niyazov et al. (2024) in different form. It is convenient to write the resulting formulas, using the parameters $\lambda,$ $R_{\theta}$ and $R_{\phi}$ instead of the parameters $t,$ $\theta$ and $\phi$, according to the following definitions:
$t=e^{-\lambda},\quad r=\sqrt{1-e^{-2\lambda}},\quad 0<\lambda<\infty,$ (56)
$R_{\theta}=\sin^{2}\theta,$ (57)
$R_{\phi}=\sin^{2}(2\pi\phi).$ (58)
Direct calculation of energy averages in formulas (14) and (15) yields
${\mathcal{T}}^{\mathrm{HI}}=C\sum_{m,n=0}^{1}A_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,$, (59)
${\mathcal{T}}_{2}^{\mathrm{HI}}=\tilde{C}\sum_{m,n=0}^{3}\tilde{A}_{(m,n)}\frac{\sinh^{2n}\lambda}{D^{m}}\,$, (60)
with
$D$ $=\sinh^{2}(2\lambda)+R_{\theta}+(1-R_{\theta})R_{\phi}\,$, (61)
$C$ $=\tanh\lambda\,$,
$A_{(0,0)}$ $=1-R_{\theta}/2\,$,
$A_{(0,1)}$ $=0\,$,
$A_{(1,0)}$ $=R_{\theta}(R_{\theta}+(1-R_{\theta})R_{\phi})/2\,$,
$A_{(1,1)}$ $=R_{\theta}\,$,
which leads to Eq. (40) and
$\tilde{C}$ $=\tanh\lambda/\cosh^{2}\lambda\,$, (62)
$\tilde{A}_{(0,0)}$ $=1+R_{\theta}(1-R_{\theta}/2)\,$,
$\tilde{A}_{(0,1)}$ $=2-R_{\theta}(2-R_{\theta})\,$,
$\tilde{A}_{(1,0)}$ $=R_{\theta}^{2}(R_{\theta}+R_{\phi}(1-R_{\theta}))/4\,$,
$\tilde{A}_{(1,1)}$ $=R_{\theta}(R_{\theta}-4(1-R_{\theta})(3-2R_{\theta})(1-R_{\phi}))/2\,$,
$\tilde{A}_{(2,0)}$ $=-R_{\theta}(2-R_{\theta})(R_{\theta}+R_{\phi}(1-R_{\theta}))^{2}/2\,$,
$\tilde{A}_{(2,1)}$ $=-R_{\theta}(R_{\theta}(1-R_{\phi})+R_{\phi})$
[…]
$\tilde{A}_{(3,0)}$ $=-R_{\theta}^{2}(R_{\theta}+R_{\phi}(1-R_{\theta}))^{3}/4\,$,
$\tilde{A}_{(3,1)}$ $=-3R_{\theta}^{2}(R_{\theta}+R_{\phi}(1-R_{\theta}))^{3}/2\,$,
$\tilde{A}_{(3,2)}$ $=-3R_{\theta}^{2}(R_{\theta}+R_{\phi}(1-R_{\theta}))^{2}\,$,
$\tilde{A}_{(3,3)}$ $=-R_{\theta}^{2}(4-2R_{\theta}(1-R_{\phi})-2R_{\phi})$
$\times(R_{\theta}(1-R_{\phi})+R_{\phi})\,$,
$\tilde{A}_{(0,2)}$ $=\tilde{A}_{(0,3)}=\tilde{A}_{(1,2)}=\tilde{A}_{(1,3)}=\tilde{A}_{(2,3)}=0\,.$
C.2 Non-identical contacts
In general case the tunneling rates in different contacts of the interferometer are described by unequal tunneling amplitudes
$t_{\mathrm{R,L}}$ $=te^{\pm\beta}\,,\quad|\beta|\leq\lambda$ (63)
$t_{\mathrm{R}}t_{\mathrm{L}}$ $=e^{-2\lambda}\,$,
$t_{\mathrm{R}}/t_{\mathrm{L}}$ $=e^{2\beta}\,.$
In this case we may still use the expressions (59), (41), but the coefficients should be redefined, $A_{(m,n)}\to A_{(m,n)}^{\mathrm{u}}$, $\tilde{A}_{(m,n)}\to\tilde{A}_{(m,n)}^{\mathrm{u}}$, as follows
$C^{\mathrm{u}}=$ $C\left(1-\frac{\sinh^{2}\beta}{\sinh^{2}\lambda}\right)\,$, (64)
$A_{(0,0)}^{\mathrm{u}}=$ $A_{(0,0)}\,$,
$A_{(0,1)}^{\mathrm{u}}=$ $0\,$,
$A_{(1,0)}^{\mathrm{u}}=$ $A_{(1,0)}+R_{\theta}\sinh^{2}\beta\,$,
$A_{(1,1)}^{\mathrm{u}}=$ $A_{(1,1)}+2R_{\theta}\sinh^{2}\beta\,$,
$\tilde{C}^{\mathrm{u}}=$ $\tilde{C}\left(1-\frac{\sinh^{2}\beta}{\sinh^{2}\lambda}\right)^{2}\,$,
$\tilde{A}_{(0,0)}^{\mathrm{u}}=$ $\tilde{A}_{(0,0)}+2R_{\theta}(2-R_{\theta})\sinh^{2}\beta\,$,
$\tilde{A}_{(0,1)}^{\mathrm{u}}=$ $\tilde{A}_{(0,1)}\,$,
[…]
$-3(1-R_{\theta})(4-3R_{\theta})R_{\phi}+4R_{\theta}\sinh^{2}\beta)\sinh^{2}\beta/2\,$,
$\tilde{A}_{(1,1)}^{\mathrm{u}}=$ $\tilde{A}_{(1,1)}+R_{\theta}^{2}\sinh^{2}(2\beta)\,$,
$\tilde{A}_{(2,0)}^{\mathrm{u}}=$ $\tilde{A}_{(2,0)}+R_{\theta}(R_{\theta}(1-R_{\phi})+R_{\phi})$
[…]
$-4(2-R_{\phi}))\sinh^{2}\beta/2$
$+5R_{\theta}^{2}(1-R_{\theta})(1-R_{\phi})\sinh^{4}\beta\,$,
$\tilde{A}_{(2,1)}^{\mathrm{u}}=$ $\tilde{A}_{(2,1)}-4R_{\theta}(2-R_{\theta})(R_{\theta}(1-R_{\phi})+R_{\phi})\sinh^{2}\beta$
[…]
$+4R_{\phi})\sinh^{2}\beta\,$,
$\tilde{A}_{(3,0)}^{\mathrm{u}}=$ $\tilde{A}_{(3,0)}-R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})$
$\times((R_{\theta}(1-R_{\phi})+R_{\phi})(2-R_{\theta}(1-R_{\phi})-R_{\phi})$
$+(4-3R_{\theta}(1-R_{\phi})-3R_{\phi})\sinh^{2}\beta)\sinh^{2}\beta\,$,
$\tilde{A}_{(3,1)}^{\mathrm{u}}=$ $\tilde{A}_{(3,1)}-6R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})(R_{\theta}(1-R_{\phi})+R_{\phi}$
$+(2-R_{\theta}(1-R_{\phi})-R_{\phi})\sinh^{2}\beta)\sinh^{2}\beta\,$,
[…]
$+R_{\phi}+\sinh^{2}\beta)\sinh^{2}\beta\,$,
$\tilde{A}_{(3,3)}^{\mathrm{u}}=$ $\tilde{A}_{(3,3)}-2R_{\theta}^{2}(R_{\theta}(1-R_{\phi})+R_{\phi})\sinh^{2}(2\beta)\,$,
$\tilde{A}_{(0,2)}^{\mathrm{u}}=$ $\tilde{A}_{(0,3)}^{\mathrm{u}}=\tilde{A}_{(1,2)}^{\mathrm{u}}=\tilde{A}_{(1,3)}^{\mathrm{u}}=\tilde{A}_{(2,3)}^{\mathrm{u}}=0\,.$